# Resource Allocation via Backscatter-Aware Transmit Antenna Selection for Low-PAPR and Ultra-Reliable WSNs

### v1

@@ V-J Joint Optimization and Fundamental Trade-off
As illustrated in Fig. 9, the system exhibits a clear operating boundary with a distinct ”knee” that defines the limits of selection diversity. The smoothed boundary (Savitzky-Golay filtered) reveals the optimal operating ”knee” where significant suppression is achieved with minimal BER penalty. By increasing the victim weight $\lambda_{V}$, the system achieves a maximum interference suppression of approximately $6.5$ dB. The ”Optimal Operating Point” in the figure is identified at $\Delta I\approx 4.5$ dB, where the system achieves a significant $2.8\times$ reduction in linear interference power while maintaining a primary link BER on the order of $10^{-3}$.
This point represents the threshold of maximum marginal utility. In the region prior to this point ($0\leq\Delta I\leq 4.5$ dB), the BC-TAS algorithm successfully identifies antenna-subcarrier mappings that are naturally favorable to the victim with negligible impact on the primary link. However, beyond this $4.5$ dB threshold, the system enters a regime of diminishing returns; further gains in interference mitigation force the selection of antennas with significantly lower channel gains for the primary receiver, resulting in an exponential increase in BER as the diversity gain is exhausted to satisfy the strict coexistence constraint.
@@
The impact of hardware complexity on this trade-off is further explored in Fig. 10. As $N_{tx}$ increases, the Pareto front shifts the entire operating boundary toward the high-suppression, low-BER region.. This shift indicates that higher-order selection diversity provides a more favorable trade-off surface, allowing for simultaneous improvements in both reliability and coexistence. For $N_{tx}=8$, the system can achieve suppression levels exceeding $7$ dB while remaining well below the BER levels of a $N_{tx}=2$ system with zero suppression.
V-K Spatial Correlation, PAPR, and Backscatter Efficiency
Conventional multi-user MIMO (MU-MIMO) systems and Space-Time Block Coding (STBC) schemes face significant challenges in WSN deployments, particularly regarding PAPR and backscatter energy harvesting efficiency. While linear receivers like MMSE detectors experience severe noise enhancement under high transmit-side spatial correlation ($\mathbf{R}_{\mathrm{tx}}$), STBC requires multiple active RF chains, which increases hardware cost and limits the power available for backscatter modulation.
In contrast, the proposed BC-TAS scheme addresses these bottlenecks by exploiting selection diversity across spatial and frequency domains. By activating only a single transmit antenna per subcarrier, BC-TAS inherently solves the high PAPR issue typically found in multi-stream MIMO, as the localized power concentration reduces the dynamic range of the transmitted signal. Furthermore, by identifying the most favorable propagation paths that maximize backscatter reflection power while minimizing interference, the scheme ensures high energy efficiency for passive sensor nodes.
@@ V-K Spatial Correlation, PAPR, and Backscatter Efficiency
As illustrated in Fig. 11, at a fixed SNR of $16$ dB, the MMSE baseline exhibits an exponential rise in BER, reaching approximately $2.8\times 10^{-2}$ at $\mathbf{R}_{\mathrm{tx}}=0.9$. While STBC provides better stability through orthogonal diversity, it maintains a performance gap significantly wider than the proposed scheme.
The proposed BC-TAS demonstrates superior robustness, achieving a BER of approximately $4.5\times 10^{-4}$ at extreme correlation levels ($0.9$). This result confirms that BC-TAS effectively circumvents the ill-conditioned channel matrices that plague MMSE and the hardware overhead of STBC. By focusing transmit energy into a single optimal antenna-subcarrier mapping, BC-TAS maximizes the signal-to-noise ratio for both the primary receiver and the backscatter tag, providing a low-complexity solution for high-density, correlated WSN environments.
V-L Physical Layer Insights: Frequency-Selective Notching
To provide physical intuition into the BC-TAS mechanism, we analyze the Spectral Power Density (SPD) of the transmitted signal as observed at the victim receiver in a frequency-selective TGn Model F environment.
@@ V-L Physical Layer Insights: Frequency-Selective Notching
As illustrated in Fig. 12, when the system operates in ”Protective Mode” ($\lambda_{V}=10$), the algorithm selectively maps subcarriers to antennas that exhibit deep fading at the victim’s spatial location. This results in a ”notching” effect where the interference power is surgically removed from specific subcarriers. For $\lambda_{V}=10$, the BC-TAS algorithm identifies and exploits natural channel nulls across the subcarrier index to protect the victim.
Crucially, this observed ”notching” validates the spatial-frequency decoupling inherent in our algorithm. It proves that the interference nulling is achieved through channel-aware selection diversity rather than expensive digital filtering or active cancellation hardware—computational overheads that a passive backscatter tag cannot support. By exploiting the natural multipath environment to ”hide” the signal from the victim, the BC-TAS framework justifies its low hardware complexity while maintaining high primary link power on subcarriers where the victim is naturally shielded.
VI Discussion and Theoretical Validation
In this section, we synthesize our numerical findings with the analytical framework derived in Section III and Appendix A. The alignment between the observed gains and the order-statistics model provides a robust validation of the BC-TAS framework.
VI-A Linking Field Flattening to Lemma 1
The results in Fig. 4 provide empirical confirmation of the ”Field Flattening” effect described in Lemma 1. The $1.1$ dB BCF advantage over OES at $N_{t}=32$ demonstrates that selection diversity can be dual-purposed for both link reliability and peak power suppression.
The use of 1,000 Monte Carlo trials confirms that this is not a transient benefit but a persistent statistical property of the BC-TAS heuristic. By smoothing the measurement noise through the Kalman update (Section III-C), the system effectively ”regularizes” the cost function, ensuring that the selected antenna-subcarrier mapping represents a stable propagation path rather than a noisy outlier. This stability is critical for low-complexity backscatter tags that lack the processing power to track high-variance reflected waveforms.
VI-B Linking Interference Suppression to SDR Expansion
The observed $2.4$ dB PAPR reduction and the corresponding interference suppression ”knee” in Fig. 9 directly validate the result in Lemma 1. As established analytically, the selection of the minimum interference path among $N_{t}=4$ antennas yields an expected power reduction scaling by $1/N_{t}$.
In the logarithmic domain, this theoretical shift is approximately $10\log_{10}(4)\approx 6.02$ dB. Our empirical maximum suppression of $6.5$ dB (see Fig. 9) aligns closely with this bound. The expansion of the SDR is therefore a direct consequence of this ”PDF squeezing” effect, where the background interference floor $\mathbb{E}[P_{B}|\lambda_{T}]$ is lowered, allowing the SAW sensor to operate in a significantly cleaner spectral environment.
VI-C Validation of the three-order-of-magnitude Outage Reduction
The massive reliability improvement reported in Table III is mathematically consistent with the diversity order $L=N_{t}$. Theoretically, the outage probability for selection diversity follows $P_{out}\approx(P_{out,SISO})^{N_{t}}$.
Given our SISO baseline outage of $0.105$ (at $16$ dB), the theoretical prediction for $N_{t}=4$ is $(0.105)^{4}\approx 1.2\times 10^{-4}$. Our simulated BC-TAS outage probability of $1.75\times 10^{-4}$ tracks this theoretical floor with high precision. The slight delta is attributed to the ”Selection Penalty” $\Delta$, where the system occasionally sacrifices the absolute maximum SNR to satisfy the weight constraints $\lambda_{T}$ and $\lambda_{V}$, as derived in Appendix A.C.
VI-D The PAPR-Linearity-Efficiency Loop
The $2.4$ dB gain in the PAPR CCDF (Fig. 6) serves as the hardware-level justification for the EVM and Energy Efficiency gains. By selecting antenna mappings that truncate the high-power tail of the OFDM distribution, BC-TAS forces the signal into the linear region of the Rapp PA model.
This enables the transmitter to reduce the IBO by approximately $3$ dB while maintaining 16-QAM compliance (see Fig. 5). This ”liberated” power is redirected toward tag illumination, increasing the harvesting efficiency $\eta_{H}$. Consequently, the system closes a performance loop: selection diversity reduces PAPR, which improves PA efficiency, which ultimately expands the power budget for the passive backscatter link.
@@ A-A Standard TAS Baseline
$\mathbb{E}[G_{\max}]=\sigma_{L}^{2}\sum_{n=1}^{N_{t}}\frac{1}{n}\approx\sigma_{L}^{2}\ln(N_{t})+\gamma_{e}$ (13)

### v2

@@ VI-J Joint Optimization and Fundamental Trade-off
As illustrated in Fig. 9, the system exhibits a clear operating boundary with a distinct “knee” defining the limits of selection diversity. By increasing the victim weight $\lambda_{V}$, maximum suppression reaches $\approx 6.5$ dB. The optimal operating point is identified at $\Delta I\approx 4.5$ dB, where the system achieves a $2.8\times$ reduction in linear interference power while maintaining a primary link BER of $10^{-3}$. In the region $\Delta I\leq 4.5$ dB, BC-TAS identifies antenna-subcarrier mappings naturally favorable to the victim with negligible BER penalty. Beyond this threshold, the system enters a regime of diminishing returns where further interference mitigation forces the selection of significantly weaker primary channel gains, resulting in an exponential BER increase. The empirical 6.5 dB maximum suppression aligns with the $10\log_{10}(N_{t})\approx 6.02$ dB theoretical bound, validating the “PDF squeezing” effect. This confirms that the SDR expansion is a direct consequence of lowering the interference floor $\mathbb{E}[P_{B}|\lambda_{T}]$, allowing the SAW sensor to operate in a cleaner spectral environment.
@@
The impact of hardware complexity is explored in Fig. 10. As $N_{tx}$ increases, the Pareto front shifts toward the high-suppression, low-BER region, indicating that higher-order selection diversity provides a more favorable trade-off surface. For $N_{tx}=8$, the system achieves suppression exceeding $7$ dB while remaining significantly more reliable than an $N_{tx}=2$ system with zero suppression.
VI-K Spatial Correlation, PAPR, and Backscatter Efficiency
Conventional MU-MIMO and Space-Time Block Coding (STBC) schemes face challenges in WSNs regarding PAPR and energy harvesting efficiency. While MMSE detectors suffer noise enhancement under high spatial correlation ($\mathbf{R}_{\mathrm{tx}}$) and STBC requires multiple costly RF chains, BC-TAS exploits selection diversity to bypass these bottlenecks. By activating a single antenna per subcarrier, BC-TAS reduces signal dynamic range, inherently solving the multi-stream PAPR issue while maximizing backscatter reflection power.
@@ VI-K Spatial Correlation, PAPR, and Backscatter Efficiency
As shown in Fig. 11, at 16 dB SNR, the MMSE baseline exhibits an exponential BER rise, reaching $2.8\times 10^{-2}$ at $\mathbf{R}_{\mathrm{tx}}=0.9$. Although STBC offers better stability, BC-TAS demonstrates superior robustness, achieving a BER of $\approx 4.5\times 10^{-4}$ at the same extreme correlation. This confirms that BC-TAS circumvents the ill-conditioned matrices inherent in MMSE and the hardware overhead of STBC. By focusing transmit energy into optimal antenna-subcarrier mappings, BC-TAS maximizes the SNR for both the primary receiver and the backscatter tag, providing a low-complexity solution for high-density, correlated WSN environments.
VI-L Physical Layer Insights: Frequency-Selective Notching
To provide physical intuition, we analyze the Spectral Power Density (SPD) at the victim receiver in a frequency-selective TGn Model F environment.
@@ VI-L Physical Layer Insights: Frequency-Selective Notching
As illustrated in Fig. 12, operating in ”Protective Mode” ($\lambda_{V}=10$) forces the algorithm to map subcarriers to antennas exhibiting deep fading at the victim’s location. This creates a ”notching” effect where interference power is surgically removed from specific subcarriers by exploiting natural channel nulls. Crucially, this validates the spatial-frequency decoupling inherent in BC-TAS: interference nulling is achieved through channel-aware selection rather than expensive digital filtering or active cancellation hardware. By utilizing the multipath environment to ”hide” the signal from the victim, BC-TAS maintains low hardware complexity while concentrating transmit power on subcarriers where the victim is naturally shielded.
@@ A-A Standard TAS Baseline
$BC-TAS\mathbb{E}[G_{\max}]=\sigma_{L}^{2}\sum_{n=1}^{N_{t}}\frac{1}{n}\approx\sigma_{L}^{2}\ln(N_{t})+\gamma_{e}$ (13)